# RBI-ThPhys-2025-26

### v2

@@ Acknowledgements.
Appendix A Theory Expressions for Tree-level Tensor Bispectrum
@@
$\mathcal{K}_{ij}^{\delta,(1)}(\mathbf{k})$ $=\frac{1}{3}\delta_{ij};$ (58)
$\mathcal{K}_{ij}^{\text{g},(1)}(\mathbf{k})$ $=b_{1}^{\text{g}}\big(\frac{\mathbf{k}_{i}\mathbf{k}_{j}}{k^{2}}-\frac{1}{3}\delta_{ij}\big)=b_{1}^{\text{g}}\mathcal{D}_{ij}(\mathbf{k});$
where in the second line we introduced the trace-free Fourier space operator $\mathcal{D}_{ij}(\mathbf{k})=\mathbf{k}_{i}\mathbf{k}_{j}/k^{2}-\frac{1}{3}\delta_{ij}$. The second order kernels are given by
$\mathcal{K}_{ij}^{\delta,(2)}(\mathbf{k}_{1},\mathbf{k}_{2})$ $=\frac{1}{3}\delta_{ij}F_{2}(\mathbf{k}_{1},\mathbf{k}_{2});$ (59)
@@
$\Delta_{ij,kl}^{\text{TF}}=\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{3}\delta_{ij}\delta_{kl}$ (66)
@@
Appendix B Details on Covariance and Model Fits
@@ B.1 Further comparison between Gaussian and sample covariance
Figure 14: Ratio of estimated and Gaussian variance for the three parity-odd bispectra for which we predict a nonvanishing signal. The grey shade indicates the region where $k_{\text{max}}<0.05\,h/\text{Mpc}$; deviations in this region are likely due to discretization effects and do not indicate model failure.
@@ B.2 Undoing chi-squared bias when using the diagonal sample covariance
However, when $\hat{C}$ is the sample covariance, the second term in Eq. (72) is independent of the data and simply reads
@@
$\hat{\sigma}_{\text{red}}\equiv(\text{Var}(\hat{\chi}_{\text{red}}^{2}))^{1/2}=\bigg(\text{Var}(\frac{\hat{\bar{\chi}}_{*}^{2}}{N_{T}N_{L}})\bigg)^{1/2}=\bigg(\frac{2(N_{r}-1)^{2}(N_{r}-2)}{N_{r}^{2}N_{T}N_{L}(N_{r}-N_{L}-2)^{2}(N_{r}-N_{L}-4)}\bigg)^{1/2}.$ (79)
@@
That is, the scatter in the reduced chi-squared computed from summing all the individual realizations is smaller when using the sample covariance than when using a noiseless covariance. To our knowledge, this result has not been appreciated in the literature thus far. On the other hand, if one does not sum the individual realizations but rather uses the mean data vector (with the appropriately rescaled covariance), the above calculation shows that the expectation of $\sigma_{\text{red}}\sim\sqrt{2/(N_{T}N_{L})}$ is correctly recovered both when using the sample covariance and when using a noiseless covariance.
@@
Appendix C Line-of-sight dependence of Shape Field
@@
Figure 15 shows the ratio of measurements of the $(\ell_{1},\ell_{2})=(0,0)$ and $(\ell_{1},\ell_{2})=(2,0)$ moments of the collinear bispectrum. We also consider the ratio of the lowest order multipole to the other quadrupole, that is $(\ell_{1},\ell_{2})=(1,1)$. For these wavenumber configurations, we see that the ratios are roughly scale-independent and in accordance with their theoretically predicted values (given the discretization effects we already encountered in the main text, we do expect to see some deviations at the level shown in the figure). We thus conclude that this discrepancy is not a source of systematics for our purposes.

### v3

@@ Acknowledgements.
Appendix A A: Theory Expressions for Tree-level Tensor Bispectrum
@@
$\mathcal{K}_{ij}^{\delta,(1)}(\mathbf{k})$ $=\frac{1}{3}\delta_{ij}^{\mathrm{K}};$ (58)
$\mathcal{K}_{ij}^{\text{g},(1)}(\mathbf{k})$ $=b_{1}^{\text{g}}\big(\frac{\mathbf{k}_{i}\mathbf{k}_{j}}{k^{2}}-\frac{1}{3}\delta_{ij}^{\mathrm{K}}\big)=b_{1}^{\text{g}}\mathcal{D}_{ij}(\mathbf{k});$
where in the second line we introduced the trace-free Fourier space operator $\mathcal{D}_{ij}(\mathbf{k})=\mathbf{k}_{i}\mathbf{k}_{j}/k^{2}-\frac{1}{3}\delta_{ij}^{\mathrm{K}}$. The second order kernels are given by
$\mathcal{K}_{ij}^{\delta,(2)}(\mathbf{k}_{1},\mathbf{k}_{2})$ $=\frac{1}{3}\delta_{ij}^{\mathrm{K}}F_{2}(\mathbf{k}_{1},\mathbf{k}_{2});$ (59)
@@
$\Delta_{ij,kl}^{\text{TF}}=\delta_{ik}^{\mathrm{K}}\delta_{jl}^{\mathrm{K}}+\delta_{il}^{\mathrm{K}}\delta_{jk}^{\mathrm{K}}-\frac{2}{3}\delta_{ij}^{\mathrm{K}}\delta_{kl}^{\mathrm{K}}$ (66)
@@
Appendix B B: Details on Covariance and Model Fits
@@ B.1 Further comparison between Gaussian and sample covariance
Figure 14: Ratio of estimated and Gaussian variance for the three parity-odd bispectra for which we predict a nonvanishing signal. The grey shade indicates the region where $k_{\text{max}}<0.05\,h/\text{Mpc}$; deviations in this region are likely due to discreteness effects and do not indicate model failure.
@@ B.2 Undoing chi-squared bias when using the diagonal sample covariance
Note that Eq. (72) is true regardless of the structure of the covariance matrix in question. However, when $\hat{C}$ is the sample covariance, the second term in Eq. (72) is independent of the data and simply reads
@@
$\hat{\sigma}_{\text{red}}\equiv(\text{Var}(\hat{\chi}_{\text{red}}^{2}))^{1/2}=\bigg(\text{Var}\bigg(\frac{\hat{\bar{\chi}}_{*}^{2}}{N_{T}N_{L}}\bigg)\bigg)^{1/2}=\bigg(\frac{2(N_{r}-1)^{2}(N_{r}-2)}{N_{r}^{2}N_{T}N_{L}(N_{r}-N_{L}-2)^{2}(N_{r}-N_{L}-4)}\bigg)^{1/2}.$ (79)
@@
That is, the scatter in the reduced chi-squared computed from summing all the individual realizations is smaller when using the sample covariance than when using a noiseless covariance. To our knowledge, this result has not been appreciated in the literature thus far. On the other hand, if one does not sum the individual realizations but rather uses the mean data vector (with the covariance appropriately rescaled by the number of realizations), the above calculation shows that the expectation of $\sigma_{\text{red}}\sim\sqrt{2/(N_{T}N_{L})}$ is correctly recovered both when using the sample covariance and when using a noiseless covariance.
@@
Appendix C C: Line-of-sight dependence of Shape Field
@@
Figure 15 shows the ratio of measurements of the $(\ell_{1},\ell_{2})=(0,0)$ and $(\ell_{1},\ell_{2})=(2,0)$ moments of the collinear bispectrum. We also consider the ratio of the lowest order multipole to the other quadrupole, that is $(\ell_{1},\ell_{2})=(1,1)$. For these wavenumber configurations, we see that the ratios are roughly scale-independent and in accordance with their theoretically predicted values (given the discreteness effects we already encountered in the main text, we do expect to see some deviations at the level shown in the figure). We thus conclude that this discrepancy is not a source of systematics for our purposes.